solve.
-3(x+1)=2(x-8)+3

Answer by checkley75(3666)

-3(x+1)=2(x-8)+3
-3x-3=2x-16+3
-3x-2x=-16+3+3
-5x=-10
x=-10/-5
x=2 answer.